without graphing, tell wether the graphsof the pairs equations are parallel,perpendicular,or neither 9x-3y=12, 27x+9y=38
Answer by mananth(16946)

9x-3y=12,
Re write the equations in slope intercept form
3y=9x-12
y=3x-4
slope m= 3

27x+9y=38
9y=-27x+38
y=-3x+38/9
the slopes are 3 & -3
if they were parallel then the slopes must be same. but the slopes are 3 & -3
If the lines are perpendicular then m1*m2=-1
3*-3 =-9
They are neither parallel nor perpendicular.
They intersect each other.